# Hidden Metallic Iron in Amorphous Silicate Dust? Insights from Condensation Experiments and Mid-Infrared Spectroscopy

### v1

@@ 3.1 Bulk analysis (XRD and EPMA)
Figure 1 shows the XRD patterns of the products from both runs and the starting material of Run CI-1. Almost all of the starting materials were vaporized during the experiments. The condensates of both runs showed broad halos in the range of , indicating the presence of amorphous materials, along with a few peaks derived from residual starting materials. Strong peaks corresponding to kamacite were observed in the product of Run CI-1. Additionally, crystalline peaks of periclase (MgO), silicon, and quartz (SiO2) were detected, which originated from the unvaporized reagents in the starting materials. Forsterite (Mg2SiO4) is supposed to be newly crystallized from the unvaporized melt, as discussed in Kim2021. The weak peak intensities of the crystalline phases indicate that the phases from unvaporized reagents are present in negligible amounts compared with the amorphous silicates.
Table 3 shows the bulk chemical composition of products determined by EPMA and the initial compositions of their starting materials. The products exhibit lower content for some elements (Mg, Ca, Na, and Al), and higher Fe and Ni contents compared to the initial compositions. This discrepancy occurred because the starting materials were not fully vaporized, as well as the surface roughness and the high porosity of the pressed samples reduced the total weight percentage measured by EPMA. The Na content was especially low (to approximately 17% of the initial value), because Na dissipated as a gas phase during the experiments due to its high volatility.
@@ 3.2 TEM observation
Figure 2 presents bright field (BF) TEM images and STEM-EDS maps of the products. The condensates were spherical amorphous silicate grains with a size of approximately 10 to in diameter. In CI-1, kamacite particles ( in diameter) were also observed. These metallic particles were supposed to be embedded inside amorphous silicate grains, as confirmed with 3D STEM-EDS tomography in the previous studies (Matsuno2021; Kim2021). The diameter ratio of a kamacite particle and whole silicate grain, , varied from 0 to 0.87 (Figure 3), and the average value was approximately 0.50. Note that not all particles had a metallic core and core size ratios were only examined for particles larger than in diameter.
The chemical compositions of amorphous grains and kamacite particles were measured with STEM-EDS. The STEM beam conditions were optimized to avoid beam damage on the amorphous silicate grains. Only maps showing no significant temporal change of textures and elemental compositions were used to derive the chemical compositions. In CI-2, the Na count was below the qualification limit. The representative chemical compositions of amorphous silicate particles were determined from the data with the highest counts. The average Mg/Si ratio of amorphous silicate and Fe/(Fe+Ni) ratio in kamacite are shown in Table 4. The Mg/Si ratios among grains were homogeneous in both runs. The relative depletion of Mg content accounts for the residue of periclase (MgO) in starting materials as seen in XRD patterns (Figure 1). In CI-1, not all Fe in the starting material condensed as kamacite but a small fraction of Fe was incorporated into the amorphous silicate structure as Fe2+, modifying the SiO4 networks of amorphous silicate with Fe/(Fe+Mg) 0.08.
@@ 3.3 Infrared spectroscopy
The IR absorbance spectra of experimental products show broad peaks at approximately and (Figure 4). These peaks are derived from Si–O stretching and Si–O–Si bending vibrations in amorphous silicates, respectively. There was no difference in positions of 9.8 and peaks between CI-1 and CI-2.
For IR reflectance measurements of polished samples such as quenched glass, a gold mirror is generally used as a reference. However, unlike polished samples, the pressed pellets of nanopowders used in this study are not as flat as the gold mirror. To account for the reflectance decrease caused by surface roughness, pressed pellets coated with thin () and thick () gold films were prepared. The thin () gold coating did not alter the original surface roughness due to the agglomerated condensed nanoparticles, while thick () coating modified the surface by filling in the the roughness (Fig. 5).
@@
In both CI-1 and CI-2 runs, the reflectance near was was approximately 0.15 larger for the thick Au-coated pellets than for the thin Au-coated pellets. This difference indicates that reflectance decreases with increasing the porosity on the pellet surface. The surface roughness was preserved on the thin Au-coated pellets better than the thick Au-coated pellets, as shown in Fig. 5. Also, the reflectance of thick and thin Au-coated pellets of CI-2 was systematically larger than those of CI-1. It is inferred that materials under the Au-coating affect the reflectance.
@@
Subsequently, we corrected the scattering effect of the thin Au-coated pellet of CI-2. The reflectance spectrum of the thin Au-coated pellet of CI-2 was approximated by a linear function of the wave number. The correction function is obtained so that the slope of the approximation line equals 1, the same as the spectrum of the gold mirror. Using the correction function, the reflectance spectra of products measured with a thin Au-coated pellet of CI-2 were modified to parallel the spectra using a gold mirror in the high wavenumber region. Fig. 7 shows the result of the correction, comparing the reflectance spectra of the iron-free product (CI-2) measured using the thin Au-coated pellet before and after the correction, and the spectrum measured using a gold mirror.
@@ 3.4 Determining optical constants of experimental products
The absorbance spectra of the products dispersed in KBr pellet are not appropriate to compare with the observed dust emission because the KBr medium effects change the peak positions of absorbance spectra (Bohren1983; Tamanai2006). The optical constants of the amorphous silicate nanoparticles condensed in the run CI-2 were determined from IR absorbance and reflectance spectra by using the Lorentz oscillator model (Bohren1983). The complex refractive index is expressed as a function of wavenumber (Equation 2).
Here, , , and represent resonance frequency, resonance strength, and damping coefficient of the -th oscillator, respectively. is dielectric constant at high wavenumbers. was determined to be 1.95 by fitting the reflectance spectra at wavenumbers above using Equation 3 as follows:
where is angle of incident light in the measurement (). We noted that the reflectance calculated using the optical constants derived from the absorbance spectrum as below did not match the measured one at the wavenumbers shorter than . This is probably due to absorption by the deformed or aggregated particles on the surface of the pressed pellet, while particles for absorbance measurements were spherical grains well dispersed in KBr. The reflectance spectra at high wavenumbers were used to determine because there is no molecular vibration resonances.
The extinction coefficient () of the product was calculated from the measured transmittance A, as follows:
Where is the average radius of the particle that satisfies as observed with TEM (Fig. 2), is the average density of amorphous silicate, M is the average mass of amorphous silicate in the pellet () and S is the surface area of the pellet (). The average density was measured with a gas displacement pycnometer (AccuPycII 1340; Shimadzu Techno-Research, Inc.) to be . In contrast to the reflectance measurement, the absorbance spectra include the KBr medium effects. Generally, there is a large uncertainty in optical constants determined from absorbance due to grain shape variety. All condensates in this study, however, are spherical shape as confirmed in TEM images. Therefore, extinction coefficient () of particles in a medium can be expressed with and as follows assuming the spherical grain shape (Koike1995);
where represents the refractive index of the medium ( in KBr). The oscillator parameters (, , and ) were changed until the calculated by Equation 5 fits the calculated from the measured absorbance spectrum in the range of as shown in Fig. 9.
The optical constants and , which were determined by fitting the absorbance spectra at using , are shown in Fig. 10 and Table LABEL:table:nk.
Finally, the absorbance spectra of spherical grains in vacuum were calculated to compare with the observations of circumstellar dust assuming , the refractive index of vacuum (Fig. 11). The peaks become weaker and shift shorter wavelengths in vacuum than in the KBr medium.
@@ 4 Discussion
4.1 Evaluation of a new method to determine optical constants of amorphous silicate nanoparticles
Previously, optical constants of amorphous silicate dust analogues have been mainly obtained by measuring the reflectance of silicate glasses synthesized with the sol-gel method (Jaeger2003) or melt-quenching (Dorschner1995; Mutschke1998; Speck2015). This approach allows the derivation of optical constants without relying on uncertain assumptions, as the reflectance spectra of glasses are straightforward to measure. In contrast, measuring the reflectance of amorphous silicate nanoparticles has been challenging due to scattering effects caused by the surface roughness of pelletized powders. Although alternative methods to derive optical constants without reflectance measurements have been proposed (Koike1989; Koike1995), was treated as a free fitting parameter or assumed (Koike1989; Koike1995; Takigawa2019). In this study, we measured both reflectance and absorbance spectra to derive the optical constants of condensed amorphous silicate nanoparticles. Ideally, optical constants derived solely from absorbance would also account for the measured reflectance and vice versa. However, as discussed in Section 3.4, the optical constants determined by independently fitting the absorbance spectra were not consistent with measured reflectance, probably because grain deformation of pressed pellets changed silicate features at . Consequently, we used the reflectance spectra at the limited range of high wavenumbers to determine and derived optical constants by fitting the absorbance spectra based on the Lorentz oscillator model. This approach is reliable to grains with known simple shapes, as the perfectly spherical grains (Fig. 2) Compared to the previous approach to deriving optical constants of non-spherical grains without reflectance measurements (Koike1989; Koike1995), our method appears more quantitative.
@@ 4.1 Evaluation of a new method to determine optical constants of amorphous silicate nanoparticles
The reflectance at high wavenumbers of our samples with a pyroxene-like Mg/Si ratio (Ca0.05Mg0.92Na0.06Al0.07Si) has a smaller value of 0.029 than the 0.045 of MgSiO3 glass made with sol-gel method (Jaeger2003) and melt-quenching (Dorschner1995). Structures such as porosity and density can differ depending on the synthesis method (Speck2011). These structural differences among silicate glasses and amorphous silicate particles may lead to the different reflectance at high wavenumbers. Different types of samples used to measure the reflectance can cause variations in the value of reflectance. While the samples in this study and Jaeger2003 are powders, the quenched glass in Dorschner1995 is the bulk material with a smooth surface.
To estimate the effects of different reflectance at high wavenumbers on the determination of and optical constants, we derived the optical constants for CI-2 (Ca0.05Mg0.92Na0.06Al0.07Si) using the reflectance at high wavenumbers reported for MgSiO3 glass (Jaeger2003; Dorschner1995) of 2.35. A comparison of derived optical constants and normalized spectra of in vacuum using and are shown in Fig. 12 and Fig. 13. The refractive index derived with was slightly lower than that determined with , while the extinction coefficient showed little difference. In normalized spectra of , the 10 and 18 m peak exhibited negligible shifts, within (from 9.42 to and from 17.8 to ) when varied from 1.95 to 2.35. These results indicate that a 36 difference in reflectance at high wavenumbers has a limited impact on spectral changes. However, further studies are necessary to refine methods for measuring the reflectance of nanoparticles and determining optical constants with higher accuracy.
@@ 4.2 Comparison of measured and modeled spectra of amorphous silicate particles with metallic iron
The STEM analysis showed that the Fe-Ni-including product with CI chondritic composition (CI-1) had a metallic core and silicate shell structure with the averaged core size ratio () of (Figs. 2 and 3). In contrast, the product in the run CI-2 was pure amorphous silicate particles. The amorphous silicate shell of CI-1 and silicate grain of CI-2 had almost the same chemical composition (Table 4). In spite of the structural difference between CI-1 and CI-2, the peak position of amorphous silicate feature did not change as the result of absorption measurement (Fig. 4).
@@
We calculated the extinction coefficient of amorphous silicates with metallic Fe-Ni cores using the optical constants of the iron-free product with the CI chondritic composition (CI-2) obtained in Section 3.4. and -iron (Ordal1985). Here, we assumed and the medium effect of KBr to compare with the measured absorption spectrum of CI-1 nanoparticles dispersed in a KBr pellet. In the calculation, we used the FORTRAN code DMiLay (Ackerman1981). This code computes electromagnetic scattering by a particle composed of a spherical core and a spherical shell without assumptions such as effective medium approximations. The measured spectrum of the product in the run CI-1 and the calculated spectrum of particles with metallic core () were compared in Fig. 14. The optical constants of metallic iron were taken from the literature (Ordal1985). There was a difference of about in the peak position between the measured and modeled spectra.
@@
This discrepancy is likely due to the assumption in the calculation that all dust grains have an identical value of , whereas the experimental products exhibit a distribution of core size ratio, as shown in Fig. 3. It should also be noted that there is a size-dependent bias in the STEM observations. In STEM-EDS analysis, only particles larger than were counted because X-ray counts obtained from smaller particles without electron beam damage were insufficient to determine the boundaries between the metallic core and silicate mantle. In addition, many small particles () were stuck to the larger grains, and their metallic cores were not confirmed in the STEM-DF images (Fig. 2). Figure 15 shows the calculated spectra of core-shell dust with various radius ratios. The peak positions do not change dramatically for grains with , whereas the peak positions of the grains with significantly shift to the shorter wavelength. The identical MIR features observed in CI-1 and CI-2, despite the presence of metallic cores in some large silicate grains, suggest that the infrared absorption spectrum is predominantly governed by numerous small particles (), for which metallic cores were not confirmed.
@@
The elemental abundance of iron in our product CI-1 (Fe/Si 0.82) was almost the CI chondritic composition (Fe/Si 0.85). Contrary to the intuitive expectation that the metallic iron of the Solar abundance would have a significant impact on the spectra, our results suggest that the spectral feature is predominantly controlled by the number density and size distribution of metallic inclusions rather than their total elemental abundance.
@@
Some CP-IDPs show IR spectral features similar to those of circumstellar and interstellar amorphous silicate dust (Bradley et al. 1999). Although the observed GEMS are aggregated and their grain boundaries are hard to distinguish, the size of metallic iron particles distributed from approximately 5 to in the silicate aggregates () (Fig. 3 of Bradley1999). The few observed isolated GEMS particles showed core-shell ratios varying from 0.35 to 0.50 (Fig. 8 of Messenger2015). Like the experimental product CI-1, metallic iron particles embedded in GEMS exhibit a size distribution. As suggested by our experimental results, metallic particles with a size distribution are difficult to detect in the mid-infrared spectra when observed in space. Therefore, circumstellar dust exhibiting MIR spectral features similar to those of GEMS may also contain metallic particles. As the presence of metallic cores enhances absorption efficiency in near-infrared wavelengths (Jaeger2003), optical constants of amorphous silicate with a metallic iron core need to be determined at a wider wavelength to further verify the presence of metallic cores in circumstellar dust.
@@
Mg-silicate condenses at a higher temperature than metallic iron because of its low surface energy under rapid cooling conditions(Yamamoto1977) . The spherical shape of the condensed grains indicates that amorphous silicate condensed as melt droplets, as proposed by previous studies (Kim2021; Matsuno2021; Enju2022). In the run CI-1, metallic iron particles condense on the surface and then move into the inside of the silicate melt (Matsuno2021; Kim2021). The iron particles in the silicate melt coagulate to grow into a large core to minimize their surface energy, leading to a heterogeneous size distribution from core-less amorphous silicate grains to large cores () in relatively large silicate grains as seen in our experiment. This can reduce the interaction cross-section of the metal particles with light and diminish their spectral contribution. In this way, the presence of metallic iron may be ”hidden” in the infrared spectra despite its high elemental abundance. Such processes could occur in rapidly cooled dense gas like the rear of shock waves in circumstellar environments or chondrule-forming regions in the protosolar disk.

### v2

@@ 3.1 Bulk analysis (XRD and EPMA)
Figure 1 shows the XRD patterns of the products of both runs and the starting material of Run CI-1. Almost all of the starting materials were vaporized during the experiments. The condensates of both runs showed broad halos in the range , indicating the presence of amorphous materials, along with a few peaks derived from residual starting materials. Strong peaks corresponding to kamacite were observed in the product of Run CI-1. Additionally, crystalline peaks of periclase (MgO), silicon, and quartz (SiO2) were detected, which originated from the unvaporized reagents in the starting materials. Forsterite (Mg2SiO4) is supposed to be newly crystallized from the unvaporized melt, as discussed in Kim2021. The weak peak intensities of the crystalline phases indicate that the phases from unvaporized reagents are present in negligible amounts compared with the amorphous silicates.
Table 3 shows the bulk chemical composition of products determined using an EPMA and the initial compositions of their starting materials. The products exhibit lower content for some elements (Mg, Ca, Na, and Al), and higher Fe and Ni contents compared to the initial compositions. This discrepancy occurred because the starting materials were not fully vaporized, as well as the surface roughness and the high porosity of the pressed samples reduced the total weight percentage measured by an EPMA. The Na content was especially low (to approximately 17% of the initial value), because Na dissipated as a gas phase during the experiments due to its high volatility.
@@ 3.2 TEM observation
Figure 2 presents bright field TEM images and STEM-EDS maps of the products. The condensates were spherical amorphous silicate grains with a size of approximately 10 to in diameter. In CI-1, kamacite particles ( in diameter) were also observed. These metallic particles were supposed to be embedded inside amorphous silicate grains, as confirmed with 3D STEM-EDS tomography in the previous studies (Matsuno2021; Kim2021). The diameter ratio of a kamacite particle and whole silicate grain, , varied from 0 to 0.87 (Fig. 3), and the average value was approximately 0.50. Note that not all particles had a metallic core and core size ratios were only examined for particles larger than in diameter.
The chemical compositions of amorphous grains and kamacite particles were measured with STEM-EDS. The STEM beam conditions were optimized to avoid beam damage on the amorphous silicate grains. Only maps showing no significant temporal change of textures and elemental compositions were used to derive the chemical compositions. In CI-2, the Na count was below the qualification limit. The representative chemical compositions of amorphous silicate particles were determined from the data with the highest counts. The average Mg/Si ratio of amorphous silicate and the Fe/(Fe+Ni) ratio in kamacite are shown in Table 4. The Mg/Si ratios were homogeneous among the grains and consistent between the two runs. The relative depletion of Mg content accounts for the residue of periclase (MgO) in starting materials as seen in XRD patterns (Fig. 1). In CI-1, not all Fe in the starting material condensed as kamacite, but a small fraction of Fe was incorporated into the amorphous silicate structure as Fe2+, modifying the SiO4 networks of amorphous silicate with Fe/(Fe+Mg) 0.08.
@@ 3.3 Infrared spectroscopy
The IR absorbance spectra of experimental products show broad peaks at approximately and (Fig. 4). These peaks are derived from Si–O stretching and Si–O–Si bending vibrations in amorphous silicates, respectively. There was no difference in the positions of the 9.8 and peaks between CI-1 and CI-2.
For IR reflectance measurements of polished samples such as quenched glass, a gold mirror is generally used as a reference. However, unlike polished samples, the pressed pellets of nanopowders used in this study are not as flat as the gold mirror. To account for the decrease in reflectance caused by surface roughness, pressed pellets coated with thin () and thick () gold films were prepared. The thin () gold coating did not alter the original surface roughness due to the agglomerated condensed nanoparticles, while the thick () coating modified the surface by filling in the roughness (Fig. 5).
@@
In both the CI-1 and CI-2 runs, the reflectance near was approximately 0.15 larger for the thick Au-coated pellets than for the thin Au-coated pellets. This difference indicates that reflectance decreases with increasing porosity on the pellet surface. The surface roughness was preserved on the thin Au-coated pellets better than the thick Au-coated pellets, as shown in Fig. 5. Also, the reflectance of thick and thin Au-coated pellets of CI-2 was systematically larger than that of CI-1. It is inferred that materials under the Au-coating affect the reflectance.
@@
Subsequently, we corrected the scattering effect of the thin Au-coated pellet of CI-2. The reflectance spectrum of the thin Au-coated pellet of CI-2 was approximated by a linear function of the wave number. The correction function is obtained so that the slope of the approximation line equals 1, the same as the spectrum of the gold mirror. Using the correction function, the reflectance spectra of products measured with a thin Au-coated pellet of CI-2 were modified to parallel the spectra using a gold mirror in the high wavenumber region. Figure 7 shows the result of the correction, comparing the reflectance spectra of the iron-free product (CI-2) measured using the thin Au-coated pellet before and after the correction, and the spectrum measured using a gold mirror.
@@ 3.4 Determining optical constants of experimental products
The absorbance spectra of the products dispersed in a KBr pellet are not appropriate to compare with the observed dust emission because the KBr medium effects change the peak positions of the absorbance spectra (Bohren1983; Tamanai2006). The optical constants of the amorphous silicate nanoparticles condensed in the run CI-2 were determined from IR absorbance and reflectance spectra by using the Lorentz oscillator model (Bohren1983). The complex refractive index is expressed as a function of wavenumber, (Eq. 2):
Here, , , and represent the resonance frequency, resonance strength, and damping coefficient of the -th oscillator, respectively. is the dielectric constant at high wavenumbers. was determined to be 1.95 by fitting the reflectance spectra at wavenumbers above using Eq. 3 as follows:
where is angle of incident light in the measurement (). We note that the reflectance calculated using the optical constants derived from the absorbance spectrum, as below, did not match the measured one at the wavenumbers shorter than . This is probably due to absorption by the deformed or aggregated particles on the surface of the pressed pellet, while particles for absorbance measurements were spherical grains well dispersed in KBr. The reflectance spectra at high wavenumbers were used to determine because there is no molecular vibration resonances.
The extinction coefficient () of the product was calculated from the measured transmittance A as follows:
Here is the average radius of the particle that satisfies as observed with TEM (Fig. 2), is the average density of amorphous silicate, M is the average mass of amorphous silicate in the pellet () and S is the surface area of the pellet (). The average density was measured with a gas displacement pycnometer (AccuPycII 1340; Shimadzu Techno-Research, Inc.) to be . In contrast to the reflectance measurement, the absorbance spectra include the KBr medium effects. Generally, there is a large uncertainty in optical constants determined from absorbance due to grain shape. All condensates in this study, however, were spherical in shape as confirmed in TEM images. Therefore, extinction coefficient () of particles in a medium can be expressed with and as follows assuming the spherical grain shape (Koike1995):
where represents the refractive index of the medium ( in KBr). The oscillator parameters (, , and ) were changed until the calculated by Eq. 5 fits the calculated from the measured absorbance spectrum in the range of as shown in Fig. 9.
The optical constants and , which were determined by fitting the absorbance spectra at using , are shown in Fig. 10 and Table LABEL:table:nk. Using these optical constants, we calculated the absorbance spectra of spherical grains in vacuum to compare with the observations of circumstellar dust assuming , the refractive index of vacuum (Fig. 11). The peaks become weaker and shift to shorter wavelengths in vacuum than in the KBr medium.
@@ 4 Discussion
4.1 Evaluation of a new method for determining optical constants of amorphous silicate nanoparticles
Previously, optical constants of amorphous silicate dust analogs have been mainly obtained by measuring the reflectance of silicate glasses synthesized with the sol-gel method (Jaeger2003) or melt-quenching (Dorschner1995; Mutschke1998; Speck2015). This approach allows the derivation of optical constants without relying on uncertain assumptions, as the reflectance spectra of glasses are straightforward to measure. In contrast, measuring the reflectance of amorphous silicate nanoparticles has been challenging due to scattering effects caused by the surface roughness of pelletized powders. Although alternative methods for deriving optical constants without reflectance measurements have been proposed (Koike1989; Koike1995), was treated as a free fitting parameter or assumed (Koike1989; Koike1995; Takigawa2019). In this study, we measured both reflectance and absorbance spectra to derive the optical constants of condensed amorphous silicate nanoparticles. Ideally, optical constants derived solely from absorbance would also account for the measured reflectance and vice versa. However, as discussed in Sect. 3.4, the optical constants determined by independently fitting the absorbance spectra were not consistent with measured reflectance, probably because grain deformation of pressed pellets changed silicate features at . Consequently, we used the reflectance spectra at the limited range of high wavenumbers to determine and derived optical constants by fitting the absorbance spectra based on the Lorentz oscillator model. This approach is reliable for grains with known simple shapes, as the perfectly spherical grains observed in this study (Fig. 2) allow the use of Mie theory without additional shape assumptions. Compared to the previous approach to deriving optical constants of nonspherical grains without reflectance measurements (Koike1989; Koike1995), our method appears more quantitative.
@@ 4.1 Evaluation of a new method for determining optical constants of amorphous silicate nanoparticles
The reflectance at high wavenumbers of our samples with a pyroxene-like Mg/Si ratio (Ca0.05Mg0.92Na0.06Al0.07Si) has a smaller value of 0.029 than the 0.045 of MgSiO3 glass made with sol-gel method (Jaeger2003) and melt-quenching (Dorschner1995). Structures such as porosity and density can differ depending on the synthesis method (Speck2011). These structural differences among silicate glasses and amorphous silicate particles may lead to different reflectance at high wavenumbers. Different types of samples used to measure the reflectance can cause variations in the value of reflectance. While the samples in this study and Jaeger2003 are powders, the quenched glass in Dorschner1995 is the bulk material with a smooth surface.
To estimate the effects of different reflectance at high wavenumbers on the determination of and optical constants, we also derived the optical constants for CI-2 (Ca0.05Mg0.92Na0.06Al0.07Si) by adopting , a value consistent with the high-wavenumber reflectance reported for MgSiO3 glass (Jaeger2003; Dorschner1995). A comparison of derived optical constants and normalized spectra of in vacuum using and is shown in Fig. 12 and Fig. 13. The refractive index derived with was slightly lower than that determined with , while the extinction coefficient showed little difference. In normalized spectra of , the 10 and 18 m peak exhibited negligible shifts, within (from 9.42 to and from 17.8 to ) when varied from 1.95 to 2.35. These results indicate that a 36 difference in reflectance at high wavenumbers has a limited impact on spectral changes. However, further studies are necessary to refine methods for measuring the reflectance of nanoparticles and determining optical constants with higher accuracy.
@@ 4.2 Comparison of measured and modeled spectra of amorphous silicate particles with metallic iron
The STEM analysis showed that the Fe-Ni-including product with CI chondritic composition (CI-1) had a metallic core and silicate shell structure with the averaged core size ratio () of (Figs. 2 and 3). In contrast, the product in run CI-2 was pure amorphous silicate particles. The amorphous silicate shell of CI-1 and the silicate grain of CI-2 had almost the same chemical composition (Table 4). Despite the structural difference between CI-1 and CI-2, the peak position of the amorphous silicate feature did not change as a result of the absorption measurement (Fig. 4).
@@
We calculated the extinction coefficient of amorphous silicates with metallic Fe-Ni cores using the optical constants of the iron-free product with the CI chondritic composition (CI-2) obtained in Sect. 3.4. and -iron (Ordal1985). Here, we assumed and the medium effect of KBr to compare with the measured absorption spectrum of CI-1 nanoparticles dispersed in a KBr pellet. In the calculation, we used the FORTRAN code DMiLay (Ackerman1981). This code computes electromagnetic scattering by a particle composed of a spherical core and a spherical shell without assumptions such as effective medium approximations. The measured spectrum of the product in the run CI-1 and the calculated spectrum of particles with metallic core () were compared in Fig. 14. The optical constants of metallic iron were taken from the literature (Ordal1985). There was a difference of about in the peak position between the measured and modeled spectra.
@@
This discrepancy is likely due to the assumption in the calculation that all dust grains have an identical value of , whereas the experimental products exhibit a distribution of core size ratio, as shown in Fig. 3. It should also be noted that there is a size-dependent bias in the STEM observations. In STEM-EDS analysis, only particles larger than were counted because X-ray counts obtained from smaller particles without electron beam damage were insufficient to determine the boundaries between the metallic core and silicate mantle. In addition, many small particles () were stuck to the larger grains, and their metallic cores were not confirmed in the STEM dark field images (Fig. 2). Figure 15 shows the calculated spectra of core-shell dust with various radius ratios. The peak positions do not change dramatically for grains with , whereas the peak positions of the grains with significantly shift to the shorter wavelength. The identical MIR features observed in CI-1 and CI-2, despite the presence of metallic cores in some large silicate grains, suggest that the IR absorption spectrum is predominantly governed by numerous small particles (), for which metallic cores were not confirmed.
@@
The elemental abundance of iron in our product CI-1 (Fe/Si of 0.82) was almost the CI chondritic composition (Fe/Si of 0.85). Contrary to the intuitive expectation that the metallic iron of the solar abundance would have a significant impact on the spectra, our results suggest that the spectral feature is predominantly controlled by the number density and size distribution of metallic inclusions rather than their total elemental abundance.
@@
Some chondritic porous interplanetary dust particles show IR spectral features similar to those of circumstellar and interstellar amorphous silicate dust (Bradley et al. 1999). Although the observed GEMS are aggregated and their grain boundaries are hard to distinguish, the size of metallic iron particles distributed from approximately 5 to in the silicate aggregates (; Fig. 3 of Bradley1999). The few observed isolated GEMS particles showed core-shell ratios varying from 0.35 to 0.50 (Fig. 8 of Messenger2015). Like the experimental product CI-1, metallic iron particles embedded in GEMS exhibit a size distribution. As suggested by our experimental results, metallic particles with a size distribution are difficult to detect in the MIR spectra when observed in space. Therefore, circumstellar dust exhibiting MIR spectral features similar to those of GEMS may also contain metallic particles. As the presence of metallic cores enhances absorption efficiency in NIR wavelengths (Jaeger2003), optical constants of amorphous silicate with a metallic iron core need to be determined at a wider wavelength to further verify the presence of metallic cores in circumstellar dust.
@@
Mg-silicate condenses at a higher temperature than metallic iron because of its low surface energy under rapid cooling conditions(Yamamoto1977). The spherical shape of the condensed grains indicates that amorphous silicate condensed as melt droplets, as proposed in previous studies (Kim2021; Matsuno2021; Enju2022). In the run CI-1, metallic iron particles condense on the surface and then move into the inside of the silicate melt (Matsuno2021; Kim2021). The iron particles in the silicate melt coagulate to grow into a large core to minimize their surface energy, leading to a heterogeneous size distribution from core-less amorphous silicate grains to large cores () in relatively large silicate grains as seen in our experiment. This can reduce the interaction cross section of the metal particles with light and diminish their spectral contribution. In this way, the presence of metallic iron may be ”hidden” in the IR spectra despite its high elemental abundance. Such processes could occur in rapidly cooled dense gas, like the rear of shock waves in circumstellar environments or chondrule-forming regions in the protosolar disk.